If marginal propensity to save is 3/4, then the marginal propensity to consume is
1
Expert's answer
2019-04-17T09:49:08-0400

MPC or Marginal propensity to consume is also given as, (1-MPS).  Where, MPS is marginal propensity to save.

So, if MPS =3/4

Then MPC = (1-3/4)

                 = 1/4 or 0.25